Circular No. 8147

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003fd IN UGC 8670 An apparent supernova in UGC 8670 has been found via unfiltered

CCDs independently by M. Papenkova and W. Li from LOTOSS/KAIT images (cf. IAUC 8143) and by T. Puckett and D. Ireland from Puckett Observatory 0.36-m reflector images (cf. IAUC 8117). The KAIT position for the new object is α = 13^{h}41^{m}43^{s}.70, δ = +40^{o}52′38′′.5 (equinox 2000.0), which is 16′′.2 east and 9′′.4 north of the nucleus of UGC 8670. Puckett provides position end figures 43^{s}.76, 37′′.2 for SN 2003fd. Available approximate magnitudes: Apr. 20, [20.0 (Puckett); May 28.2 UT, [19.5 (KAIT); June 9.2, 17.8 (KAIT); 10.175, 17.4 (Puckett); 10.2, 17.8 (KAIT).

SUPERNOVAE 2003fe, 2003ff, 2003fg, 2003fh, 2003fi, 2003fj, 2003fk P. Astier, Laboratoire de Physique Nucléaire et de Hautes Energies,

Université Paris VI; and C. Pritchet, University of Victoria; on behalf of the Canada-France-Hawaii Telescope (CFHT) Legacy Survey Supernova Project (including E. Aubourg, D. Balam, R. Carlberg, D. Hardin, D. Guide, S. Fabbro, H. Lafoux, T. Merrall, R. Pain, K. Perrett, J. Rich, K. Schahmaneche, P. Antilogus, E. Lesquoy, and G. Sainton), report the discovery of several apparent supernovae found on CCD images taken with the newly commissioned MegaPrime instrument during the CFHT Legacy Survey engineering pre-survey. The tabulated magnitudes below are in the i′ band. None of the new objects were present in reference images (limiting mag ∼ 24.5; S/N = 5 for 2′′ aperture) taken on 2003 Mar. 3 (SNe 2003fe, 2003ff, 2003fg, 2003fh) and 23 (SNe 2003fi, 2003fj, 2003fk).

SN  2003 UT  α_{2000}  δ_{2000}  i′  Offset
2003fe Mar. 23.53 14^{h}17^{m}19.^{s}12 +52^{o}14′43′′.8 23.3 0′′.4 W, 1′′.2 N
2003ff Mar. 23.53 14 21 14.92 +52 32 15.3 23.6 3′′.5 E, 2′′.8 S
2003fg Apr. 24.54 14 16 18.78 +52 14 55.4 20.6 0′′.2 W
2003fh Apr. 24.54 14 21 35.89 +52 31 37.7 23.2 0′′.4 W, 1′′.3 S
2003fi May 24.45 14 18 39.96 +52 36 44.1 23.2 0′′.6 E, 1′′.6 S
2003fj May 24.45 14 19 45.19 +52 32 25.7 22.2 1′′.0 W, 1′′.0 N
2003fk May 24.45 14 19 55.84 +53 05 50.9 22.4 1′′.8 W, 2′′.6 S

Additional i′ magnitudes: SN 2003fe, Mar. 25.5 UT, 23.0; Apr. 1.5, 22.9. 2003ff, Mar. 25.5, 23.3; Apr. 1.5, 23.1. 2003fg, Apr. 29.5, 20.4. 2003fh, Apr. 29.5, 22.2. 2003fi, May 28.4, 22.8; June 2.4, 22.4. 2003fj, May 28.4, 22.3; June 2.4, 22.5. SN 2003fk, May 28.4, 22.3; June 2.4, 22.3.

2003 June 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green